How many classes of solutions are there for each of the following congruences?

(a) x2 - 1 = 0 mod (168)

(b) x2 + 1 = 0 mod (70)

(c) x2 + x + 1 = 0 mod (91)

(d) x3 + 1 = 0 mod (140)

Please note to show how you got the solutions as well as finding out how many classes of solutions there are for each congruence.  

Please explain every step so I can understand how tonot only the solutions but the classes of solutions.   
Answer #1

23 II n = ti 168 = 34 7x8 x²_1=0 (mod 3) han sols : or x = 1, 2 (mod 3) x-1=0 brood 7) hon sochs: x = lg6 modt) x-1=0 forod 8